Solve the given initial value problem. = 4x + y - e (0) = 2 dt dy dt =x+uya (0) = -3 The solution is x(t)= and y(t)=
Answer #1

solution :- Given = 4x+y - 4t da at dy dt = 22+ 4y ; 4(0)-2,400)=-3 This can be written in Matrix form as e46 2 X= 54 where[i][b]-[: il aty zo vector corresponding is an [ . to vecters t ces Czest y=c, este When a=30 ③ becomes, il, x=-4 So eigen a=we let X is a of particular solution (Here C, and Ca are constants to be determined.) is the non past. Here [osta] shomogeneo36 ses 8 2 Thus XC+) = ciest y(t)= C, est+cest text is the general solution of Now to solve the given IUP, we use the given c